OpenStudy (anonymous):

if (u, v) = u-v/u+v, find F(1/u, 1/v) + F(u,v). the only question left to finish my assignment :)) please please, anyone. Our teacher haven't discussed this yet, and i can't find any examples in the internet..

OpenStudy (anonymous):

or any idea what is this called? :)

OpenStudy (whpalmer4):

painful algebra :-) \[F(u,v) = \frac{u-v}{u+v}\] \[F(1/u,1/v) = \frac{1/u-1/v}{1/u+1/v}\] \[F(1/u,1/v)+F(u,v) = \frac{1/u-1/v}{1/u+1/v} +\frac{u-v}{u+v}\] Grind through the algebra...

OpenStudy (anonymous):

ans just simplify the last one? :) ohhhhhhhhh, thank you. hahahahaha :) algebra algebra algebra. lol.

OpenStudy (anonymous):

and*

OpenStudy (anonymous):

"Grind through the algebra..." xD LOL you can throw it into wolfram to confirm your answer

OpenStudy (anonymous):

wolfram? is it for free? haha, like openstudy? :)

OpenStudy (whpalmer4):

Here's a little trick that makes this much easier: substitute c = 1/u, d = 1/v and do the algebra until you can't get any further, then substitute back.

OpenStudy (jhannybean):

Cross multiply :)

OpenStudy (whpalmer4):

The problem becomes \[\frac{c-d}{c+d}+\frac{u-v}{u+v}\]

OpenStudy (whpalmer4):

Make a common denominator and combine, then substitute back. I think you'll be pleasantly surprised by the result :-)

OpenStudy (anonymous):

wolfram's basic function are free. copy/paste this: (1/u - 1/v)/(1/u + 1/v) + (u-v)/(u+v) here: http://www.wolframalpha.com/ i was about to say what palmer said about the result :P

OpenStudy (jdoe0001):

http://openstudy.com/study#/updates/51bcb172e4b0286018b16d2b in case you missed on your previouis posting of this duplicate post

OpenStudy (anonymous):

ahh okayy :) i got 0 as the result.. hehe

OpenStudy (anonymous):

youre right

OpenStudy (jhannybean):

\[F(1/u,1/v)+F(u,v) = \frac{1/u-1/v}{1/u+1/v} +\frac{u-v}{u+v}\]\[\frac{1/u-1/v}{1/u+1/v} \cdot (u+v) + \frac{u+v}{u+v}\cdot (1/u+1/v)\] Yeah....you get 0 lol.

OpenStudy (anonymous):

haha thank you people! :) mwah hahaha

OpenStudy (anonymous):

@jdoe0001 oh, that's my classmate over there :) haha

OpenStudy (jdoe0001):

ahemm, yes and I noticed you didn't read it, instead just reposted it the same thing :/

OpenStudy (anonymous):

i didn't :P

OpenStudy (jdoe0001):

yes, I saw that

OpenStudy (anonymous):

i recommeded her the openstudy at first and i didnt know she posted the same thing

OpenStudy (jdoe0001):

hmmmm, I said what I said because I did read you 1st posting and gave you the link this is your 2nd posting of the same exact thing

OpenStudy (anonymous):

uhmm yeah, on mathematics,.

OpenStudy (anonymous):

the subject**